# From LLMs to LLM-based Agents for Software Engineering: A Survey of Current, Challenges and Future

### v1

@@ VII-D Benchmarks
The benchmarks include public datasets and datasets self-crafted by the authors themselves, and the application scenarios are also quite differently as shown in the Table VIII. BigCloneBench is a benchmark dataset for code clone detection, containing a large number of Java function pairs. These pairs are classified as clones and non-clones, used for training and evaluating clone detection models, with the main evaluation metric being the correct identification rate. The Chrysalis dataset created by [116], it contains over 1000 function-level designs from 11 open-source synthesizable HLS datasets, primarily used to evaluate the effectiveness of LLM debugging tools in detecting and correcting injected errors in HLS designs, with the main evaluation metric being the effectiveness of error detection and correction. The CodexGLUE dataset is a comprehensive benchmark dataset covering various code generation and understanding tasks such as code completion, code repair, and code translation, used to evaluate the performance of code generation models in practical programming tasks. In addition to these public datasets, some artificially simulated datasets are used, such as a simulated job fair environment dataset. This dataset simulates a virtual job fair environment containing multiple task scenarios such as interviews, recruitment, and team project coordination. The dataset used to evaluate the coordination capabilities of generative agents in complex social tasks, with the main evaluation metrics being task coordination success rate and role matching accuracy.
@@ VII-E Evaluation Metrics
TABLE VIII: Evaluation Metrics in Software Design and Evaluation
@@ 31 Questions from
70 User Requirements.
Number of files generated,
Time taken, Cost
@@ VIII Software Test Generation
In software development, a crucial component is software testing, which need to continuously been conducted from the initial system development to the final deployment. In industry, agile development is commonly used which test system continuously at every stage to ensure the robustness of the entire system, whenever new code is committed to the GitHub, tests are conducted to ensure the usability of the updated version. A common approach is to use Jenkins [^4] https://www.jenkins.io/ to achieve continuous integration and continuous deployment. Jenkins automatically hooks into the developer’s action of pushing code to GitHub and runs a test suite against the new version. Although the entire process leans towards automated development, creating and refining test cases still requires large human effort.
Typical roles in development involve software testing, such as writing unit tests, integration tests, and fuzz tests. Researchers have been attempting to use AI to help generate test cases since before the 2000. Initial implementations typically involved simpler forms of AI and machine learning to automate parts of the test case generation process. Over time, more sophisticated methods such as natural language processing and machine learning models have been applied to improve the precision and scope of test case generation. Online tools like Sofy [^5] https://sofy.ai/, which use machine learning to generate context-based paths in applications, also exist to aid in generating test suites. Using large language models to generate test cases is a relatively new attempt but has been developing rapidly. In 2020, researchers utilized pre-trained language models fine-tuned on labeled data to generate test cases. They developed a sequence-to-sequence transformer-based model called ”ATHENATEST” and compared its generated results with EvoSuite and GPT-3, demonstrating better test coverage [131]. More research and models are being dedicated to test suite generation experiments, for instance, the Codex model [67], mentioned earlier in the code generation section, combined with chain-of-thought prompting, achieved high-quality test suite generation with CodeCoT, even in zero-shot scenarios. The introduction of LLMs aims to automate and streamline the testing process, making it more rigorous and capable of addressing aspects that humans might easily overlook.
@@ VIII-A LLMs Tasks
The application of LLMs in software test generation is extensive and encompasses more than just test suite generation. The reviewed paper included in this survey covers several aspects, including security test generation, bug reproduction, general bug reproduction, fuzz testing, and coverage-driven test generation. These tasks are achieved through various models and techniques, significantly improving software quality and reducing developers’ workload.  [132] aims to evaluate the effectiveness of using GPT-4 to generate security tests, demonstrating how to conduct supply chain attacks by exploiting dependency vulnerabilities. The study experimented with different prompt styles and templates to explore the effectiveness of varying information inputs on test generation quality, the results showed that tests generated by ChatGPT successfully discovered 24 proof-of-concept vulnerabilities in 55 applications, outperforming existing tools TRANSFER[133] and SIEGE [^6] https://siegecyber.com.au/services/penetration-testing/. This research introduces a new method for generating security tests using LLMs and provides empirical evidence of LLM’s potential in the security testing domain, offering developers a novel approach to handling library vulnerabilities in applications.
@@
Another application is bug reproduction, which allows testers to locate and fix bugs more quickly and efficiently. [134] addresses the limitations of current bug reproduction methods, which are constrained by the quality and clarity of handcrafted patterns and predefined vocabularies. The paper proposes and evaluates a new method framework called AdbGPT, which uses a large language model to automatically reproduce errors from Android bug reports. AdbGPT is described as outperforming current SOTA approaches in the context of automated bug replay for only Android system. The experimental results show that AdbGPT achieved accuracies of 90.4% and 90.8% in S2R entity extraction and a success rate of 81.3% in error reproduction, significantly outperforming the baseline ReCDroid and ablation study versions. By introducing prompt engineering, few-shot learning, and chain-of-thought reasoning, AdbGPT demonstrates the powerful capabilities of LLMs in automated error reproduction. It also uses GUI encoding to convert the GUI view hierarchy into HTML-like syntax, providing LLMs with a clear understanding of the current GUI state. While AdbGPT is specialized for Android systems, [135] proposes the LIBRO framework, which uses LLMs to generate bug reproduction tests from bug reports. The experimental results show that LIBRO successfully reproduced 33.5% of bugs in the Defects4J dataset and 32.2% in the GHRB dataset. By combining advanced prompt engineering and post-processing techniques, LIBRO demonstrates the effectiveness and efficiency of LLMs in generating bug reproduction tests. Although LIBRO has a lower absolute effectiveness compared to AdbGPT, it was tested across a more diverse set of Java applications and not limited to Android. Therefore, while AdbGPT excels in specialized bug replay for Android, LIBRO provides a wider range of bug reproduction for Java applications. The extensive application of LLMs in test generation tasks such as security test generation, bug reproduction, fuzz testing, program repair, and coverage-driven test generation highlights their significant potential in improving software quality and reducing the burden on developers. Through various models and techniques, these tasks demonstrate how LLMs can automate and enhance the software testing process, addressing aspects that are often overlooked by humans.
@@
Similarly, in fuzz testing, LLMs have shown promise potential. [136] developed a universal fuzzing tool, Fuzz4All, which uses LLMs to generate and mutate inputs for various software systems. This tool addresses the issues of traditional fuzzers being tightly coupled with specific languages or systems and lacking support for evolving language features. The study conducted various experiments to test the tool’s capabilities, including coverage comparison, bug finding, and targeted fuzzing. The results showed that Fuzz4All achieved the highest code coverage in all tested languages, with an average increase of 36.8%, and discovered 98 bugs across nine systems, which considered as state-of-art technique in universal fuzzing with LLMs at that time. Through self-prompting and LLM-driven fuzzing loops, Fuzz4All demonstrated the effectiveness of LLMs in fuzz testing and showcased their capability across multiple languages and systems under test (SUTs) through comprehensive evaluations.
@@
[137] introduced SymPrompt, a new code-aware prompting strategy aimed at addressing the limitations of existing Search-Based Software Testing (SBST) methods and traditional LLM prompting strategies in generating high-coverage test cases. By decomposing the original test generation process into a multi-stage sequence aligned with the execution paths of the method under test, SymPrompt generated high-coverage test cases. Experimental results indicated that SymPrompt increased coverage on CodeGen2 and GPT-4 by 26% and 105% respectively. Through path constraint prompting and context construction techniques, SymPrompt demonstrated the potential of LLMs in generating high-coverage test cases. [138] also focused on test suite coverage, this study introduced the COVERUP system which generates high-coverage Python regression tests through coverage analysis and interaction with LLMs. The experimental results showed that COVERUP increased code coverage from 62% to 81% and branch coverage from 35% to 53% through iterative prompting and coverage-driven methods. [139] proposed the AID method, which combines LLMs with differential testing to improve fault detection in ”plausibly correct” software. By comparing the effectiveness of AID in generating fault-revealing test inputs and oracles, the experiments showed that AID improved recall and precision by 1.80 times and 2.65 times respectively, and increased the F1 score by 1.66 times. By integrating LLMs with differential testing, AID showcased the powerful capability of LLMs in detecting complex bugs.
@@ VIII-B LLM-based Agents Tasks
In the field of software test generation, the application of LLM-based agents demonstrates their potential in automated test generation. While relying on LLM-based agents for software test generation might seem excessive, more research is directed towards vulnerability detection and system maintenance. LLM-based agents can enhance test reliability and quality by distributing tasks such as test generation, execution, and optimization through a multi-agent collaborative system. These multi-agent systems offer obvious improvements in error detection and repair, and coverage testing. An example of such a system is AgentCoder’s multi-agent framework, as discussed in the code generation and software development section [82]. The primary goal of this system is to leverage multiple specialized agents to iteratively optimize code generation, overcoming the limitations of a single agent model in generating effective code and test cases. The paper introduce the test design agent, which creates diverse and comprehensive test cases; and the test execution agent, which executes the tests and provides feedback, it reached an 89.9% pass rate on the MBPP dataset. Similarly, the SocraTest framework falls under the Autonomous Learning and Decision Making topic [106]. This framework automates the testing process through conversational interactions, the paper presents detailed examples of generating and optimizing test cases using GPT-4, emphasizing how multi-step interactions enhance testing methods and generate test code. Experimental results show that through conversational LLMs, SocraTest can effectively generate and optimize test cases and utilize middleware to facilitate interactions between the LLM and various testing tools, achieving more advanced automated testing capabilities.
The paper collected for the software test generation topic are mostly multiple agents based system. The study [140] evaluates the effectiveness of LLMs in generating high-quality test cases and identifies their limitations. It proposes a novel multi-agent framework called TestChain. The paper evaluates StarChat, CodeLlama, GPT-3.5, and GPT-4 on the HumanEval and LeetCode-hard datasets. Experimental results show that the TestChain framework, using GPT-4, achieved 71.79% accuracy on the LeetCode-hard dataset, an improvement of 13.84% over baseline methods. On the HumanEval dataset, TestChain with GPT-4 achieved 90.24% accuracy. The TestChain framework designs agents to generate diverse test inputs, maps inputs to outputs using ReAct format dialogue chains, and interacts with the Python interpreter to obtain accurate test outputs.
@@
LLM-based agents can also be applied in user acceptance testing (UAT), [141] aims to enhance the automation of the WeChat Pay UAT process by proposing a multi-agent collaborative system named XUAT-Copilot, which uses LLMs to automatically generate test scripts. The study evaluates XUAT-Copilot’s performance on 450 test cases from the WeChat Pay UAT system, comparing it to a single-agent system and a variant without the reflection component. Experimental results show that XUAT-Copilot achieved a Pass@1 rate of 88.55%, compared to 22.65% for the single-agent system and 81.96% for the variant without the reflection component, with a Complete@1 rate of 93.03%. XUAT-Copilot employs a multi-agent collaborative framework, including action planning, state checking, and parameter selection agents, and uses advanced prompting techniques. XUAT-Copilot demonstrates the potential and feasibility of LLMs in automating UAT test script generation.
@@ VIII-C Analysis
Figure 7: Illustration of Comparison Framework Between LLM-based Agent[141] and LLM[136] in Software Test Generation
@@
A single LLM may struggle with complex, multi-step tasks. For example, in high-coverage test generation, LLMs may require more complex prompts and post-processing steps to achieve the desired results. Additionally, the quality of the generated results depends heavily on the prompt design and quality. For tasks requiring fine control and continuous optimization, a single LLM may find it challenging to deal with. As shown in Figure.LABEL:testGen, the LLM framework uses [136] as an example to demonstrate the usage of LLMs in fuzz testing, the prompt will be optimized by given code snippets (fuzz inputs), and re-select by the LLM again to choose the best prompt for the future generation. The overall framework lacks autonomy, the LLM-based agent [141] framework on the left fills this gap, as well as able to perceive the UI and interact with the skill library for the operations. The operation agent will receive any error reported by the inspection agent and do the self-reflection to refine the process autonomously. However, as previously discussed, build a LLM-based agents framework only for the software test generation task are ”overkill”, so the collected paper for LLM-based agents system generally focused on program repair by generated test cases or bug replay system, like in the Figure.LABEL:testGen, the LLM-based agent framework is actually used for automatically test the Wechat Pay system.

### v2

@@ VII-D Benchmarks
The benchmarks include public datasets and datasets self-crafted by the authors themselves, and the application scenarios are also quite differently as shown in the Table IX. BigCloneBench is a benchmark dataset for code clone detection, containing a large number of Java function pairs. These pairs are classified as clones and non-clones, used for training and evaluating clone detection models, with the main evaluation metric being the correct identification rate. The Chrysalis dataset created by [145], it contains over 1000 function-level designs from 11 open-source synthesizable HLS datasets, primarily used to evaluate the effectiveness of LLM debugging tools in detecting and correcting injected errors in HLS designs, with the main evaluation metric being the effectiveness of error detection and correction. The CodexGLUE dataset is a comprehensive benchmark dataset covering various code generation and understanding tasks such as code completion, code repair, and code translation, used to evaluate the performance of code generation models in practical programming tasks. In addition to these public datasets, some artificially simulated datasets are used, such as a simulated job fair environment dataset. This dataset simulates a virtual job fair environment containing multiple task scenarios such as interviews, recruitment, and team project coordination. The dataset used to evaluate the coordination capabilities of generative agents in complex social tasks, with the main evaluation metrics being task coordination success rate and role matching accuracy.
@@ VII-E Evaluation Metrics
TABLE IX: Evaluation Metrics in Software Design and Evaluation
@@ 31 Questions from
Seven different game environments.
Win Rate, ErrorRate
@@ VIII Software Test Generation
In software development, a crucial component is software testing, which needs to be conducted continuously from the initial system development to the final deployment. In industry, agile development is commonly used which tests the system continuously at every stage to ensure the robustness of the entire system, whenever new code is committed to GitHub, tests are conducted to ensure the usability of the updated version. A common approach is to use Jenkins [^4] https://www.jenkins.io/ to achieve continuous integration and continuous deployment. Jenkins automatically hooks into the developer’s action of pushing code to GitHub and runs a test suite against the new version. Although the entire process leans towards automated development, creating and refining test cases still requires a large human effort.
Typical roles in development involve software testing, such as writing unit tests, integration tests, and fuzz tests. Researchers have been attempting to use AI to help generate test cases since before 2000. Initial implementations typically involved simpler forms of AI and machine learning to automate parts of the test case generation process. Over time, more sophisticated methods such as natural language processing and machine learning models have been applied to improve the precision and scope of test case generation. Online tools like Sofy [^5] https://sofy.ai/, which use machine learning to generate context-based paths in applications, also exist to aid in generating test suites. Using large language models to generate test cases is a relatively new attempt but has been developed rapidly. In 2020, researchers utilized pre-trained language models fine-tuned on labeled data to generate test cases. They developed a sequence-to-sequence transformer-based model called ”ATHENATEST” and compared its generated results with EvoSuite and GPT-3, demonstrating better test coverage [158]. More research and models are being dedicated to test suite generation experiments, for instance, the Codex model [88], mentioned earlier in the code generation section, combined with chain-of-thought prompting, achieved high-quality test suite generation with CodeCoT, even in zero-shot scenarios. The introduction of LLMs aims to automate and streamline the testing process, making it more rigorous and capable of addressing aspects that humans might easily overlook.
@@ VIII-A LLMs Tasks
Security and Vulnerability Test Generation. The application of LLMs in software test generation is extensive and encompasses more than just test suite generation. The reviewed paper included in this survey covers several aspects, including security test generation, bug reproduction, general bug reproduction, fuzz testing, and coverage-driven test generation. These tasks are achieved through various models and techniques, significantly improving software quality and reducing developers’ workload.  [159] aims to evaluate the effectiveness of using GPT-4 to generate security tests, demonstrating how to conduct supply chain attacks by exploiting dependency vulnerabilities. The study experimented with different prompt styles and templates to explore the effectiveness of varying information inputs on test generation quality, the results showed that tests generated by ChatGPT successfully discovered 24 proof-of-concept vulnerabilities in 55 applications, outperforming existing tools TRANSFER[160] and SIEGE [^6] https://siegecyber.com.au/services/penetration-testing/. This research introduces a new method for generating security tests using LLMs and provides empirical evidence of LLM’s potential in the security testing domain, offering developers a novel approach to handling library vulnerabilities in applications. [97] introduces TICODER, an interactive framework that leverages test generation to clarify user intent and improve code accuracy, this work exemplifies the synergy between software testing and code synthesis, offering a novel path toward human-in-the-loop development with LLMs.
@@
Automated Bug Reproduction from Natural Language Reports. Another application is bug reproduction, which allows testers to locate and fix bugs more quickly and efficiently. [161] addresses the limitations of current bug reproduction methods, which are constrained by the quality and clarity of handcrafted patterns and predefined vocabularies. The paper proposes and evaluates a new method framework called AdbGPT, which uses a large language model to automatically reproduce errors from Android bug reports. AdbGPT is described as outperforming current SOTA approaches in the context of automated bug replay for only Android system. The experimental results show that AdbGPT achieved accuracies of 90.4% and 90.8% in S2R entity extraction and a success rate of 81.3% in error reproduction, significantly outperforming the baseline ReCDroid and ablation study versions. By introducing prompt engineering, few-shot learning, and chain-of-thought reasoning, AdbGPT demonstrates the powerful capabilities of LLMs in automated error reproduction. It also uses GUI encoding to convert the GUI view hierarchy into HTML-like syntax, providing LLMs with a clear understanding of the current GUI state. While AdbGPT is specialized for Android systems, [162] proposes the LIBRO framework, which uses LLMs to generate bug reproduction tests from bug reports. The experimental results show that LIBRO successfully reproduced 33.5% of bugs in the Defects4J dataset and 32.2% in the GHRB dataset. By combining advanced prompt engineering and post-processing techniques, LIBRO demonstrates the effectiveness and efficiency of LLMs in generating bug reproduction tests. Although LIBRO has a lower absolute effectiveness compared to AdbGPT, it was tested across a more diverse set of Java applications and not limited to Android. Therefore, while AdbGPT excels in specialized bug replay for Android, LIBRO provides a wider range of bug reproduction for Java applications. The extensive application of LLMs in test generation tasks such as security test generation, bug reproduction, fuzz testing, program repair, and coverage-driven test generation highlights their significant potential in improving software quality and reducing the burden on developers. Through various models and techniques, these tasks demonstrate how LLMs can automate and enhance the software testing process, addressing aspects that are often overlooked by humans.
@@
Universal and Language-Agnostic Fuzz Testing. Similarly, in fuzz testing, LLMs have shown promise potential. [163] developed a universal fuzzing tool, Fuzz4All, which uses LLMs to generate and mutate inputs for various software systems. This tool addresses the issues of traditional fuzzers being tightly coupled with specific languages or systems and lacking support for evolving language features. The study conducted various experiments to test the tool’s capabilities, including coverage comparison, bug finding, and targeted fuzzing. The results showed that Fuzz4All achieved the highest code coverage in all tested languages, with an average increase of 36.8%, and discovered 98 bugs across nine systems, which considered as state-of-art technique in universal fuzzing with LLMs at that time. Through self-prompting and LLM-driven fuzzing loops, Fuzz4All demonstrated the effectiveness of LLMs in fuzz testing and showcased their capability across multiple languages and systems under test (SUTs) through comprehensive evaluations.
@@
Coverage-Driven Test Generation.  [164] introduced SymPrompt, a new code-aware prompting strategy aimed at addressing the limitations of existing Search-Based Software Testing (SBST) methods and traditional LLM prompting strategies in generating high-coverage test cases. By decomposing the original test generation process into a multi-stage sequence aligned with the execution paths of the method under test, SymPrompt generated high-coverage test cases. Experimental results indicated that SymPrompt increased coverage on CodeGen2 and GPT-4 by 26% and 105% respectively. Through path constraint prompting and context construction techniques, SymPrompt demonstrated the potential of LLMs in generating high-coverage test cases. [165] also focused on test suite coverage, this study introduced the COVERUP system which generates high-coverage Python regression tests through coverage analysis and interaction with LLMs. The experimental results showed that COVERUP increased code coverage from 62% to 81% and branch coverage from 35% to 53% through iterative prompting and coverage-driven methods. [166] proposed the AID method, which combines LLMs with differential testing to improve fault detection in ”plausibly correct” software. By comparing the effectiveness of AID in generating fault-revealing test inputs and oracles, the experiments showed that AID improved recall and precision by 1.80 times and 2.65 times respectively, and increased the F1 score by 1.66 times. By integrating LLMs with differential testing, AID showcased the powerful capability of LLMs in detecting complex bugs. Beyond proposing new prompting strategies, [167] conducts a comprehensive comparison between ChatGPT and EvoSuite [^7] https://github.com/EvoSuite/evosuite(a SOTA SBST tool) in generating unit test suites. The evaluation spans correctness, readability, code coverage, and bug detection. ChatGPT achieved 55.4% average statement coverage across 207 Java classes, while EvoSuite reached 74.2%, outperforming ChatGPT in most cases. Nonetheless, ChatGPT showed advantages in generating diverse test data, semantically rich call chains, and context-aware inputs, outperforming EvoSuite in 17.9% of the classes. The study suggests that while LLMs may lack feedback mechanisms, they offer strong generalization and usability, making them suitable as lightweight, accessible test generation tools and potential complements to SBST.
@@ VIII-B LLM-based Agents Tasks
Collaborative Multi-Agent Systems for Test Case Generation. In the field of software test generation, the application of LLM-based agents demonstrates their potential in automated test generation. While relying on LLM-based agents for software test generation might seem excessive, more research is directed towards vulnerability detection and system maintenance. LLM-based agents can enhance test reliability and quality by distributing tasks such as test generation, execution, and optimization through a multi-agent collaborative system. These multi-agent systems offer obvious improvements in error detection and repair, and coverage testing. An example of such a system is AgentCoder’s multi-agent framework, as discussed in the code generation and software development section [107]. The primary goal of this system is to leverage multiple specialized agents to iteratively optimize code generation, overcoming the limitations of a single agent model in generating effective code and test cases. The paper introduce the test design agent, which creates diverse and comprehensive test cases; and the test execution agent, which executes the tests and provides feedback, it reached an 89.9% pass rate on the MBPP dataset. Similarly, the paper collected for the software test generation topic are mostly multiple agents based system. The study [168] evaluates the effectiveness of LLMs in generating high-quality test cases and identifies their limitations. It proposes a novel multi-agent framework called TestChain. The paper evaluates StarChat, CodeLlama, GPT-3.5, and GPT-4 on the HumanEval and LeetCode-hard datasets. Experimental results show that the TestChain framework, using GPT-4, achieved 71.79% accuracy on the LeetCode-hard dataset, an improvement of 13.84% over baseline methods. On the HumanEval dataset, TestChain with GPT-4 achieved 90.24% accuracy. The TestChain framework designs agents to generate diverse test inputs, maps inputs to outputs using ReAct format dialogue chains, and interacts with the Python interpreter to obtain accurate test outputs.
Interactive and Conversational Test Generation. The SocraTest framework falls under the Autonomous Learning and Decision Making topic [135]. This framework automates the testing process through conversational interactions, the paper presents detailed examples of generating and optimizing test cases using GPT-4, emphasizing how multi-step interactions enhance testing methods and generate test code. Experimental results show that through conversational LLMs, SocraTest can effectively generate and optimize test cases and utilize middleware to facilitate interactions between the LLM and various testing tools, achieving more advanced automated testing capabilities.
@@
Integrated Testing in Software Development Pipelines.  [108] proposes AGILECODER, a multi-agent framework that integrates Agile methodology into LLM-based software development which introduced in section V-B. The Tester agent generates test suites and structured testing plans based on code changes, enabling efficient bug detection and feedback. Evaluated on HumanEval, MBPP, and a new ProjectDev benchmark, AGILECODER outperforms ChatDev and MetaGPT in both pass@1 and executability showcasing the synergy between collaborative development and automated testing in LLM-based systems. LLM-based agents can also be applied in user acceptance testing (UAT), [169] aims to enhance the automation of the WeChat Pay UAT process by proposing a multi-agent collaborative system named XUAT-Copilot, which uses LLMs to automatically generate test scripts. The study evaluates XUAT-Copilot’s performance on 450 test cases from the WeChat Pay UAT system, comparing it to a single-agent system and a variant without the reflection component. Experimental results show that XUAT-Copilot achieved a Pass@1 rate of 88.55%, compared to 22.65% for the single-agent system and 81.96% for the variant without the reflection component, with a Complete@1 rate of 93.03%. XUAT-Copilot employs a multi-agent collaborative framework, including action planning, state checking, and parameter selection agents, and uses advanced prompting techniques. XUAT-Copilot demonstrates the potential and feasibility of LLMs in automating UAT test script generation.
@@ VIII-C Analysis
Figure 8: Illustration of Comparison Framework Between LLM-based Agent[169] and LLM[163] in Software Test Generation.
@@
A single LLM may struggle with complex, multi-step tasks. For example, in high-coverage test generation, LLMs may require more complex prompts and post-processing steps to achieve the desired results. Additionally, the quality of the generated results depends heavily on the prompt design and quality. For tasks requiring fine control and continuous optimization, a single LLM may find it challenging to deal with. As shown in Figure. 8, the LLM framework uses [163] as an example to demonstrate the usage of LLMs in fuzz testing, the prompt will be optimized by given code snippets (fuzz inputs), and re-select by the LLM again to choose the best prompt for the future generation. The overall framework lacks autonomy, the LLM-based agent [169] framework on the left fills this gap, as well as able to perceive the UI and interact with the skill library for the operations. The operation agent will receive any error reported by the inspection agent and do the self-reflection to refine the process autonomously. However, as previously discussed, build a LLM-based agents framework only for the software test generation task are ”overkill”, so the collected paper for LLM-based agents system generally focused on program repair by generated test cases or bug replay system, the LLM-based agent framework is actually used for automatically test the Wechat Pay system.